OpenStudy (anonymous):

A Cab company charges a fixed flag rate of 20$ and 1.40 for every mile covered. If M represents the number of miles driven , write and expression for the total cab fare as a function of the miles driven. Then solve for the total fare after the cab traveled for 36 miles.

OpenStudy (anonymous):

20+1.4m=total money 20+1.4(36)=20+50.4=70.4